A study conducted by a TV station in a particular town showed the number of televisions
per household and the corresponding probabilities for each.

Number of televisions X 1 2 3 4
Probability P(X) 0.38 0.61 0 0.03

a. What is the probability that there are 3 televisions in a household?
b. What is the probability of having 2 or more televisions in a household?
1
Expert's answer
2020-09-07T18:15:03-0400

the number of televisions per household and the corresponding probabilities for each.

Number of televisions (X ) Probability P(X)

1 0.38

2 0.61

3 0

4 0.03


a) Now

the probability that there are 3 televisions in a household = P(3)

= 0 ANSWER


b) the probability of having 2 or more televisions in a household = P(X"\geq" 2)

=P(2)+P(3)+P(4)

=0.61+0+0.03

=0.64 ANSWER